# Milky Way-est: Cosmological Zoom-in Simulations with Large Magellanic Cloud and Gaia–Sausage–Enceladus Analogs

### v1

@@ 2.3 Milky Way-est Sample
We find that 33 halos satisfy the above criteria, out of 4455 halos in the MW mass range in c125-1024; the percentage of MW-mass halos in c125-1024 that pass each cut is summarized in Table 1. We resimulate 25 of these hosts with Lagrangian volumes smaller than $\approx 1550~{}(\mathrm{Mpc}~{}h^{-1})^{3}$, for which zoom-in simulations are computationally feasible. In five of these resimulations, the LMC analog’s orbit noticeably differs from the parent box (potentially due to the butterfly effect; Genel et al. 2019), so we do not analyze these hosts. The resulting 20 zoom-ins constitute the Milky Way-est suite. [^3] One of these hosts (Halo327) was also in the Symphony MW sample (Mao et al., 2015; Nadler et al., 2023b); when comparing the two suites, we only include it in Milky Way-est. Figure 1 shows the projected dark matter density within each host halo’s virial radius at $z\approx 0$, arranged by host mass and concentration; each LMC analog halo is visible as a prominent substructure near its host’s center.
@@
Given the ±$\pm 20~{}{\rm kpc}$ range allowed for our LMC analogs’ $z=0$ distances, we analyze each Milky Way-est system at the scale factor when its LMC analog is closest to a distance of $50~{}{\rm kpc}$ from the host center, denoted $a_{\mathrm{LMC,50}}$; if there are multiple snapshots at which the LMC analog’s distance is smaller than $50~{}{\rm kpc}$, we choose the earliest time. In general, our LMC analogs’ orbits do not exactly match those in the parent box. Thus, to ensure that all LMC analogs reach an appropriate distance, some Milky Way-est simulations are analyzed slightly after $z=0$. Because our LMC analogs are constrained to accrete recently, most of them are near or at first pericenter when they reach a distance of $\approx 50~{}{\rm kpc}$, with the exception of two LMC analogs on second infall (see Section 3.2 for details). The resulting Milky Way-est analysis snapshots are all within $\approx 100~{}{\rm Myr}$ of $z=0$, and are listed in Appendix A. We refer to these as the $z\approx 0$ snapshots, and we normalize scale factors for each Milky Way-est host relative to $a_{\mathrm{LMC,50}}$. Note that we always analyze Symphony MW hosts at $z=0$.
@@ 3.1 MW Analog Characteristics
After performing our zoom-in simulations, we identified host halos by matching their accretion histories to the MAHs of target hosts in c125-1024. Our hosts’ MAHs match their counterparts in the parent box; for example, the average difference in $M_{\mathrm{host}}$ at $z\approx 0$ between our zoom-ins and the parent box is $3\%$. In our zoom-ins, all but one of the Milky Way-est hosts satisfy our MW halo concentration cut (Section 2.2), as shown by the orange circles in Figure 2. Thus, these host halo properties are robust to the change in resolution from c125-1024 to our zoom-in resimulations.
We also show Symphony MW halos (gray squares in Figure 2), which span a narrower range of host mass and a broader range of concentration than our observationally motivated cuts. When comparing Milky Way-est to Symphony MW in Section 4, we normalize subhalo masses by their corresponding host masses, which controls for the difference in the host mass distributions because subhalo abundances scale approximately linearly with host mass for MW-mass halos (e.g., Nadler et al. 2023b). At fixed host halo mass, subhalo abundances are also sensitive to host concentration (e.g., Mao et al. 2015) and other secondary properties (e.g., Fielder et al. 2019), which in turn may be affected by the presence of LMC analogs in Milky Way-est hosts (e.g., Fielder et al. 2020).
Figure 3: Maximum circular velocity, $V_{\mathrm{max,sub}}$, versus most recent infall time for Milky Way-est LMC analogs (purple circles) and the highest-$V_{\mathrm{peak,sub}}$ subhalo in each Symphony MW host (gray squares). The latest infall time is defined as the snapshot that a subhalo most recently accretes into its host. The box corresponds to our LMC analog selection cuts (i.e., criteria (IIIa) and (IIIb) in Section 2.2); Milky Way-est hosts are analyzed at the $z\approx 0$ snapshot when each LMC analog is nearest to a distance of $50~{}{\rm kpc}$ from the host center, consistent with selection criterion (IIIc).
@@ 3.2 LMC Analog Characteristics
Figure 4: Orbits of the LMC analogs in our Milky Way-est simulations (purple solid). Black filled points mark the $z\approx 0$ snapshot for each host used in our analyses, black unfilled points mark $a_{\mathrm{infall}}$, and line colors correspond to $a_{\mathrm{infall}}$. Horizontal dotted line marks $50~{}{\rm kpc}$. Additionally, solutions for the observed LMC’s orbit based on proper motion (PM) measurements are shown in black, from Patel et al. (2020); their fiducial LMC with a lower-mass (higher-mass) MW host is shown in dashed (dot-dashed). Note that all LMC analogs are on first infall, except for Halo229 and Halo659, which respectively have first pericenters at distances of $240~{}{\rm kpc}$ and $120~{}{\rm kpc}$ at $a=0.62$ and $a=0.59$.
We identify LMC analogs by matching the orbits and MAHs of subhalos in our zoom-in simulations with c125-1024. All resulting LMC analogs match the halos in the parent box well; for example, their peak masses agree at the percent level. Figure 3 shows the maximum circular velocity and accretion scale factor of each Milky Way-est LMC analog (purple circles). All Milky Way-est LMC analogs are most recently accreted within the last $1.3~{}{\rm Gyr}$ ($a_{\mathrm{infall}}\geq 0.86$, consistent with our selection criterion from Section 2.2), and almost all of these LMC analogs are on first infall. Furthermore, all of our LMC analogs have $V_{\mathrm{max,sub}}$ and $a_{\mathrm{infall}}$ values that satisfy our criteria from Section 2.2 at the percent level. At the snapshots when our LMC analogs are closest to $50~{}{\rm kpc}$ from their respective hosts’ centers, their distance distribution has a mean and host-to-host standard deviation of $r_{\mathrm{LMC}}=53\pm 11~{}{\rm kpc}$. Milky Way-est LMC analogs have a mean $V_{\mathrm{max,sub}}$ ($M_{\mathrm{sub}}$) of $94~{}{\rm km\,s^{-1}}$ ($1.8\times 10^{11}~{}{M_{\odot}}$), and are thus consistent with measurements of the LMC’s current dynamical mass (e.g., Erkal et al. 2019; Shipp et al. 2021).
@@
Figure 4 shows the orbits of all Milky Way-est LMC analogs. As noted above, all but two of our LMC analogs are near their first pericenter at the $z\approx 0$ snapshot; after infall, their orbits are generally consistent with those inferred from proper motion measurements (Kallivayalil et al., 2013), including recent studies using Gaia data (Vasiliev, 2023). We note that the orbits of our two LMC analogs on their second pericenter are also potentially consistent with orbits reconstructed from current LMC proper motion measurements (e.g., see Vasiliev 2024).
@@ 3.3 GSE Analog Characteristics
Figure 5: Scale factor at which a disrupted subhalo merges with the host, $a_{\mathrm{disrupt}}$, versus the sub-to-host halo mass ratio for this object evaluated when its peak mass is achieved, $[M_{\mathrm{sub}}/M_{\mathrm{host}}](a_{\mathrm{peak,sub}})$. Points show Milky Way-est GSE analogs (green circles) and highest-$M_{\mathrm{peak,sub}}$ Symphony MW disrupted subhalos with $0.60.25<a_{\mathrm{disrupt}}<$ 0.6 (gray squares). The box corresponds to our GSE analog selection cuts (i.e., criterion (IV) in Section 2.2).
Figure 5 shows the disruption times and sub-to-host halo mass ratios for our Milky Way-est GSE analogs (green circles). We compare these to the most massive disrupted subhalo over the relevant range of $a_{\mathrm{disrupt}}$ in each Symphony MW simulation (gray squares). We evaluate sub-to-host halo mass ratios at the snapshot when the GSE analog (or highest-$M_{\mathrm{peak,sub}}$ disrupted subhalo) achieves its peak mass to match our GSE selection criteria. For all Milky Way-est hosts except the case noted above, the GSE analog is consistent with our constraints, shown by the dashed region in Figure 5.
Our GSE analogs merge with a mean $a_{\mathrm{disrupt}}$ of $0.40\pm 0.06$, where uncertainties represent the host-to-host standard deviation. This distribution is consistent with the Symphony MW highest-$M_{\mathrm{peak,sub}}$ disrupted subhalos we compare to, which were selected over the same $a_{\mathrm{disrupt}}$ range. The mean peak mass of Milky Way-est hosts’ GSE analogs is $1.1\times 10^{11}~{}{M_{\odot}}$, versus $1.5\times 10^{11}~{}{M_{\odot}}$ for Symphony MW’s highest-$M_{\mathrm{peak,sub}}$ disrupted subhalos. Despite these similar mass distributions, many Symphony MW disrupted subhalos have sub-to-host halo mass ratios below our GSE criterion. This result is influenced by our LMC constraint—in particular, the LMC contributes significantly to Milky Way-est hosts’ total halo masses at late times, such that Milky Way-est hosts have lower masses than average during the epoch of the GSE merger (see Section 4.1). Thus, we measure larger sub-to-host halo merger mass ratios for Milky Way-est GSE analogs compared to Symphony MW highest-$M_{\mathrm{peak,sub}}$ disrupted subhalos, on average.
@@ 4.1 Host Halo Formation Histories
Figure 6: Left: Mean MAHs of Milky Way-est (gold) and Symphony MW (black solid) zoom-in hosts. The mean Symphony MW MAH normalized to the Milky Way-est value at $a=0.8$ is also shown (black dashed). Shaded regions show the 16-84th MAH percentiles for each suite. Each host’s mass is normalized to its final value before stacking ($z\approx 0$ for Milky Way-est, and $z=0$ for Symphony MW). Scale factors are normalized to the scale factor of each host’s final snapshot before averaging. Right: Mean MAHs of halos in c125-1024 that satisfy successive Milky Way-est selection criteria (see Section 2.2 and Table 1). The gray band corresponds to MW-mass halos, based on our Milky Way-est mass selection criterion, and colored lines show samples that satisfy successive cuts on host concentration (light blue long-dashed), LMC analog mass, infall time, and present-day distance (lightest to darkest purple short-dashed), and GSE analogs (gold dot-dashed).
@@
Before interpreting these results, we describe a few technical aspects of our MAH calculation. First, we present MAHs using host virial masses reported by Rockstar. This includes the masses of all subhalos within the virial radius at a given time, although we note that the late-time growth described above for Milky Way-est hosts persists even if subhalos’ masses are subtracted from the host mass at each snapshot. Second, a small subset of Milky Way-est systems have LMC analogs that are particularly massive (e.g. $M_{\mathrm{sub}}>1.5\times 10^{11}M_{\odot}$ at $z\approx 0$). However, the shift towards later assembly for Milky Way-est hosts persists even when these systems are excluded. Finally, we have checked that additional differences in host selection criteria do not drive this MAH shift. In particular, host samples from c125-1024 selected with and without the “secondary” Symphony MW host selection cuts on environment and Lagrangian volume (Nadler et al., 2023b) display consistent MAHs, implying that these characteristics do not drive the MAH difference.
@@
Figure 7: Left: Mean SHMFs, measured using present-day subhalo virial mass ($M_{\mathrm{sub}}$) and normalized by host virial mass ($M_{\mathrm{host}}$), for all subhalos within the virial radius of Milky Way-est (solid purple) and Symphony MW (solid black) hosts. Dashed lines show the contribution to these mass functions from subhalos associated with the LMC analogs in Milky Way-est and the highest-$V_{\mathrm{peak,sub}}$ subhalos in Symphony MW. Right: Mean subhalo radial distributions in Milky Way-est and Symphony MW systems as a function of distance from the host halo center, in units of $R_{\mathrm{vir,host}}$. Dashed lines show the radial distributions of subhalos associated with the LMC analog or highest-$V_{\mathrm{peak,sub}}$ subhalo, as in the left panel. In both panels, shaded bands show 16-84th percentiles, and lower panels show the ratio of Milky Way-est to Symphony MW measurements for all subhalos (solid) and non-LMC-associated subhalos (dot-dashed).
@@
We also compare Milky Way-est and Symphony MW host MAHs up to 1) $a=0.25$ (before our GSE analogs accrete) and 2) $a=0.8$ (before our LMC analogs accrete; see the black dashed line in Figure 6) to examine how these constraints affect Milky Way-est hosts’ MAHs at other times. We find that hosts’ specific accretion rates are similar between the suites for $a<0.25$. At $a=0.25$, Milky Way-est hosts are $39\%$ less massive, on average, with a mean mass of $1.0\times 10^{11}~{}{M_{\odot}}$ versus $1.6\times 10^{11}~{}{M_{\odot}}$ for Symphony MW. During GSE infall and disruption ($0.60.25<a<$ 0.6), Milky Way-est hosts’ specific accretion rates are higher than in Symphony MW, although Milky Way-est hosts are still less massive at these times, on average. At later times ($a>0.8$), Milky Way-est hosts overtake Symphony MW hosts in terms of both specific and absolute growth.

### v2

@@ 2.3 Milky Way-est Sample
We find that 33 halos satisfy the above criteria, out of 4455 halos in the MW mass range in c125-1024; the percentage of MW-mass halos in c125-1024 that pass each cut is summarized in Table 1. We then identify hosts for which zoom-in simulations are computationally feasible based on the volume of their high-resolution regions: There are 25 such hosts, and we conduct zoom-in simulations for these. In five of these resimulations, the LMC analog’s orbit noticeably differs from the parent box (potentially due to the butterfly effect; Genel et al. 2019), so we do not analyze these hosts further. The resulting 20 zoom ins constitute the Milky Way-est suite. [^3] One of these hosts (Halo327) was also in the Symphony MW sample (Mao et al., 2015; Nadler et al., 2023b); when comparing the two suites, we only include it in Milky Way-est. Figure 1 shows the projected dark matter density within each host halo’s virial radius at $z\approx 0$, arranged by host mass and concentration; each LMC analog halo is visible as a prominent substructure near its host’s center.
@@
Given the ±$\pm 20~{}{\rm kpc}$ range allowed for our LMC analogs’ $z=0$ distances, we define the snapshot at which to analyze each Milky Way-est system, $a_{\mathrm{LMC,50}}$, as the time that its LMC analog is closest to $50~{}{\rm kpc}$ from the host center. [^4] In one case (Halo659), we select the LMC’s pericenter snapshot because pericenter occurs when the LMC reaches a distance of $44~{}{\rm kpc}$, which is sufficiently close to $50~{}{\rm kpc}$ for our analysis. If there are multiple snapshots at which the LMC analog’s distance is smaller than $50~{}{\rm kpc}$, we choose the earliest time. Because our LMC analogs are constrained to accrete recently, most LMC analogs reach their first pericenter at or near this point, with the exception of two LMC analogs on second infall (see Section 3.2 for details). In general, our LMC analogs’ orbits do not exactly match those in the parent box; in order to ensure that all LMC analogs reach an appropriate distance, some Milky Way-est simulations are analyzed slightly after $z=0$. The resulting Milky Way-est analysis snapshots are typically within $\approx 500~{}{\rm Myr}$ of $z=0$, and are listed in Appendix A. We refer to these as the $z\approx 0$ snapshots, and we normalize scale factors for each Milky Way-est host relative to $a_{\mathrm{LMC,50}}$.
Although we do not explicitly constrain hosts’ accretion histories at intermediate times, nearly half of the hosts in our sample accrete a subhalo with a peak mass above $10^{11}~{}{M_{\odot}}$ that merges between GSE disruption and LMC infall, potentially analogous to Sgr; we leave a detailed study of these systems to future work. The other half of our hosts are thus quiescent at intermediate times. We have also explored SMC analogs in Milky Way-est. However, no halos that satisfy all cuts above have a clear SMC analog, i.e., a massive subhalo on a long-lived orbit around the LMC. Finally, we note that Milky Way-est halos are not explicitly selected to reside in environments representative of the Local Group or Local Volume; as a result, massive neighboring host halos reminiscent of M31 surround a handful of Milky Way-est systems. Studying the impact of Milky Way-est hosts’ environments on their accretion histories and subhalo populations is therefore an interesting area for future work.
2.4 Symphony MW Sample
Here, we briefly describe the details of the Symphony MW suite, which we use in comparative analyses with the Milky Way-est suite presented in this work.
We use $44$ Symphony MW hosts from Nadler et al. (2023b), i.e., the original sample of $45$ hosts minus Halo327; see footnote 3). These systems were first presented in Mao et al. (2015) and were selected from c125-1024 to have a mass within $10^{12.09\pm 0.02}~{}{M_{\odot}}$ at $z=0$ in the parent box. These hosts are subject to an isolation cut in the parent box, such that they cannot be within a distance of $4R_{\mathrm{vir,host}}$ from any more massive halo. Milky Way-est hosts are not subject to an explicit isolation cut. We observe in Figure 2 that Symphony MW hosts have a narrow mass distribution (due to the Symphony MW selection criteria) and wider concentration distribution than Milky Way-est hosts (due to our explicit concentration cut in Milky Way-est; criterion (II) from Table 1). To mitigate the differences in mass distributions, our subhalo population analyses are performed with subhalo masses normalized to the host mass to help isolate the effects of Milky Way-est selection criteria. Throughout this work, Symphony MW hosts are always analyzed at $z=0$.
@@ 3.1 MW Analog Characteristics
After performing our zoom-in simulations, we identified host halos by matching their accretion histories to the MAHs of target hosts in c125-1024. Our hosts’ MAHs match their counterparts in the parent box; for example, the average difference in $M_{\mathrm{host}}$ at $z\approx 0$ between our zoom ins and the parent box is $3\%$. In our zoom ins, all but one of the Milky Way-est hosts satisfy our MW halo concentration cut (Section 2.2), as shown by the orange circles in Figure 2. Thus, these host halo properties are robust to the change in resolution from c125-1024 to our zoom-in resimulations.
We also show Symphony MW halos (gray squares in Figure 2), which span a narrower range of host mass and a broader range of concentration than the Milky Way-est host sample, given our observationally motivated cuts. When comparing Milky Way-est to Symphony MW in Section 4, we normalize subhalo masses by their corresponding host masses, which controls for the difference in the host mass distributions because subhalo abundances scale approximately linearly with host mass for MW-mass halos (e.g., Nadler et al. 2023b). At fixed host halo mass, subhalo abundances are also sensitive to host concentration (e.g., Mao et al. 2015) and other secondary properties (e.g., Fielder et al. 2019), which in turn may be affected by the presence of LMC analogs in Milky Way-est hosts (e.g., Fielder et al. 2020).
@@ 3.2 LMC Analog Characteristics
We identify LMC analogs by matching the orbits and MAHs of subhalos in our zoom-in simulations with c125-1024. All resulting LMC analogs match the halos in the parent box well; for example, their peak masses agree at the percent level. Figure 3 shows the maximum circular velocity and accretion scale factor of each Milky Way-est LMC analog (purple circles). All Milky Way-est LMC analogs are most recently accreted within the last $2~{}{\rm Gyr}$ ($a_{\mathrm{infall}}\geq 0.86$, consistent with our selection criterion from Section 2.2), and almost all of these LMC analogs are on first infall. Furthermore, all of our LMC analogs have $V_{\mathrm{max,sub}}$ and $a_{\mathrm{infall}}$ values that satisfy our criteria from Section 2.2 at the percent level. At the snapshots when our LMC analogs are closest to $50~{}{\rm kpc}$ from their respective hosts’ centers, their distance distribution has a mean and host-to-host standard deviation of $r_{\mathrm{LMC}}=52\pm 11~{}{\rm kpc}$. Milky Way-est LMC analogs have a mean $V_{\mathrm{max,sub}}$ ($M_{\mathrm{sub}}$) of $94~{}{\rm km\,s^{-1}}$ ($1.8\times 10^{11}~{}{M_{\odot}}$) with a standard deviation of $22~{}{\rm km\,s^{-1}}$ ($1.4\times 10^{11}~{}{M_{\odot}}$). Of our 20 LMC analogs, 13 have present-day virial masses $M_{\mathrm{sub}}>10^{11}~{}{M_{\odot}}$ and are thus broadly consistent with measurements of the LMC’s current dynamical mass (e.g., Erkal et al. 2019; Shipp et al. 2021). We note that our LMC analog sample includes seven systems with $M_{\mathrm{sub}}<10^{11}~{}{M_{\odot}}$ because we only impose a $V_{\mathrm{max,sub}}$ cut (criterion IIIa from from Section 2.2; see Table 2 for a list of our LMC analogs’ masses).
@@
Figure 4 shows the orbits of all Milky Way-est LMC analogs. As noted above, all but two of our LMC analogs are near their first pericenter at the $z\approx 0$ snapshot; after infall, their orbits are generally consistent with those inferred from proper-motion measurements (Kallivayalil et al., 2013), including recent studies using Gaia data (Vasiliev, 2023). We note that the orbits of our two LMC analogs on their second pericenter are also potentially consistent with orbits reconstructed from current LMC proper-motion measurements (e.g., see Vasiliev 2024).
Figure 3: Maximum circular velocity, $V_{\mathrm{max,sub}}$, vs. most recent infall time for Milky Way-est LMC analogs (purple circles) and the highest-$V_{\mathrm{peak,sub}}$ subhalo in each Symphony MW host (gray squares). The latest infall time is defined as the snapshot that a subhalo most recently accretes into its host. The box corresponds to our LMC analog selection cuts (i.e., criteria (IIIa) and (IIIb) in Section 2.2); Milky Way-est hosts are analyzed at the $z\approx 0$ snapshot when each LMC analog is nearest to a distance of $50~{}{\rm kpc}$ from the host center, consistent with selection criterion (IIIc).
Figure 4: Orbits of the LMC analogs in our Milky Way-est simulations (purple solid). Black filled points mark the $z\approx 0$ snapshot for each host used in our analyses, black unfilled points mark $a_{\mathrm{infall}}$, and line colors correspond to $a_{\mathrm{infall}}$. The horizontal dotted line marks $50~{}{\rm kpc}$. Additionally, solutions for the observed LMC’s orbit based on proper-motion (PM) measurements are shown in black, from Patel et al. (2020); their fiducial LMC with a lower-mass (higher-mass) MW host is shown in dashed (dotted–dashed). Note that all LMC analogs are on first infall, except for Halo229 and Halo659, which have first pericenters at distances of $142~{}{\rm kpc}$ and $70~{}{\rm kpc}$, respectively, both at $a=0.59$.
@@ 3.3 GSE Analog Characteristics
Figure 5: Redshift at which a disrupted subhalo merges with the host, $z_{\mathrm{disrupt}}$, vs. the sub-to-host halo mass ratio for this object evaluated when its peak mass is achieved, $[M_{\mathrm{sub}}/M_{\mathrm{host}}](z_{\mathrm{peak,sub}})$. Points show Milky Way-est GSE analogs (green circles) and highest-$M_{\mathrm{peak,sub}}$ Symphony MW disrupted subhalos with $30.67<z_{\mathrm{disrupt}}<$ 3 (gray squares). The box corresponds to our GSE analog selection cuts (i.e., criterion (IV) in Section 2.2).
Figure 5 shows the disruption times and sub-to-host halo mass ratios for our Milky Way-est GSE analogs (green circles). We compare these to the most massive disrupted subhalo over the relevant range of $z_{\mathrm{disrupt}}$ in each Symphony MW simulation (gray squares). We evaluate sub-to-host halo mass ratios at the snapshot when the GSE analog (or highest-$M_{\mathrm{peak,sub}}$ disrupted subhalo) achieves its peak mass to match our GSE selection criteria. For all Milky Way-est hosts except the case noted above, the GSE analog is consistent with our constraints, shown by the dashed region in Figure 5.
Our GSE analogs merge with a mean $z_{\mathrm{disrupt}}$ of $1.54\pm 0.43$ (or mean $a_{\mathrm{disrupt}}$ of $0.40\pm 0.06$), where uncertainties represent the host-to-host standard deviation. This distribution is consistent with the Symphony MW highest-$M_{\mathrm{peak,sub}}$ disrupted subhalos we compare to, which were selected over the same disruption time range. The mean peak mass of Milky Way-est hosts’ GSE analogs is $1.1\times 10^{11}~{}{M_{\odot}}$, versus $1.5\times 10^{11}~{}{M_{\odot}}$ for Symphony MW’s highest-$M_{\mathrm{peak,sub}}$ disrupted subhalos. Despite these similar mass distributions, many Symphony MW disrupted subhalos have sub-to-host halo mass ratios below our GSE criterion. This result is influenced by our LMC constraint—in particular, the LMC contributes significantly to Milky Way-est hosts’ total halo masses at late times, such that Milky Way-est hosts have lower masses than average during the epoch of the GSE merger (see Section 4.1). Thus, we measure larger sub-to-host halo merger mass ratios for Milky Way-est GSE analogs compared to Symphony MW highest-$M_{\mathrm{peak,sub}}$ disrupted subhalos, on average.
@@ 4.1 Host Halo Formation Histories
Figure 6: Left panel: mean MAHs of Milky Way-est (gold) and Symphony MW (black solid) zoom-in hosts. The mean Symphony MW MAH normalized to the Milky Way-est value at $a=0.8$ is also shown (black dashed). Shaded regions show the 16th–84th MAH percentiles for each suite. Each host’s mass is normalized to its $z=0$ value before stacking. Scale factors are normalized to the scale factor of each host’s final snapshot before averaging. Right panel: mean MAHs of halos in c125-1024 that satisfy successive Milky Way-est selection criteria (see Section 2.2 and Table 1). The gray band corresponds to MW-mass halos, based on our Milky Way-est mass selection criterion, and colored lines show samples that satisfy successive cuts on host concentration (light blue long-dashed), LMC analog mass, infall time, and present-day distance (lightest to darkest purple short-dashed), and GSE analogs (gold dotted–dashed).
@@
Before interpreting these results, we describe a few technical aspects of our MAH calculation. First, we present MAHs using host virial masses reported by Rockstar. This includes the masses of all subhalos within the virial radius at a given time, although we note that the late-time growth described above for Milky Way-est hosts persists even if subhalos’ masses are subtracted from the host mass at each snapshot. Second, a small subset of Milky Way-est systems have LMC analogs that are particularly massive (e.g. $M_{\mathrm{sub}}>1.5\times 10^{11}M_{\odot}$ at $z\approx 0$). However, the shift toward later assembly for Milky Way-est hosts persists even when these systems are excluded. Finally, we have checked that additional differences in host selection criteria do not drive this MAH shift. In particular, host samples from c125-1024 selected with and without the “secondary” Symphony MW host selection cuts on environment and Lagrangian volume (Nadler et al., 2023b) display consistent MAHs, implying that these characteristics do not drive the MAH difference.
@@
Figure 7: Left panel: mean SHMFs, measured using present-day subhalo virial mass ($M_{\mathrm{sub}}$) and normalized by host virial mass ($M_{\mathrm{host}}$), for all subhalos within the virial radius of Milky Way-est (solid purple) and Symphony MW (solid black) hosts. Dashed lines show the contribution to these mass functions from subhalos associated with the LMC analogs in Milky Way-est and the highest-$V_{\mathrm{peak,sub}}$ subhalos in Symphony MW. Right panel: mean subhalo radial distributions in Milky Way-est and Symphony MW systems as a function of distance from the host halo center, in units of $R_{\mathrm{vir,host}}$. Dashed lines show the radial distributions of subhalos associated with the LMC analog or highest-$V_{\mathrm{peak,sub}}$ subhalo, as in the left panel. In both panels, shaded bands show 16th–84th percentiles, and lower panels show the ratio of Milky Way-est to Symphony MW measurements for all subhalos (solid) and non-LMC-associated subhalos (dotted–dashed), with black dotted lines marking a ratio of 1.
We also compare Milky Way-est and Symphony MW host MAHs up to (i) $a=0.25$ ($z=3$; before our GSE analogs accrete) and (ii) $a=0.8$ ($z=0.25$; before our LMC analogs accrete; see the black dashed line in Figure 6) to examine how these constraints affect Milky Way-est hosts’ MAHs at other times. We find that hosts’ specific accretion rates are similar between the suites for $a<0.25$. At $a=0.25$, Milky Way-est hosts are $39\%$ less massive, on average, with a mean mass of $1.0\times 10^{11}~{}{M_{\odot}}$ versus $1.6\times 10^{11}~{}{M_{\odot}}$ for Symphony MW. During GSE infall and disruption ($0.60.25<a_{\mathrm{disrupt}}<$ 0.6 or $30.67<z_{\mathrm{disrupt}}<$ 3), Milky Way-est hosts’ specific accretion rates are higher than in Symphony MW, although Milky Way-est hosts are still less massive at these times, on average. At later times ($a>0.8$), and particularly during LMC infall, Milky Way-est hosts overtake Symphony MW hosts in terms of both specific and absolute growth.